# Uncertainty in Graph Neural Networks: A Survey

### v1

@@ 4.3 Uncertainty-aware Gnn Modeling
Node-level. In graphs, nodes often exhibit conflicting or contradictory characteristics, such as belonging to different communities or revealing contradictory underlying patterns (bojchevski2017deep, ). This discrepancy should be reflected in the uncertainty of node embeddings, highlighting the need to account for uncertainty at the node-level in Gnn structures. hajiramezanali2019variational incorporate stochastic latent variables by combining a Graph Recurrent Neural Network (GRNN) with a Variational Graph Autoencoder (VGAE) (kipf2016variational, ). The proposed variational GRNN is designed to capture the temporal dependencies in dynamic graphs. Moreover, it aims to represent each node by modeling its distribution in the latent space rather than assigning it a deterministic vector in a low-dimensional space, thus enhancing its ability to represent the uncertainty associated with node latent representations. vashishth2019confidence propose ConfGCN, which introduces confidence (inversely proportional to uncertainty) estimation for the label scores on nodes to enhance label predictions. More specifically, ConfGCN assumes a fixed uncertainty for input labels and defines an influence score of each node as a parameter that can be optimized jointly with the network parameters through back-propagation. The influence score here is the inverse of the co-variance matrix-based symmetric Mahalanobis distance between two nodes in the graph. sun2021hyperbolic generate stochastic representations using a hyperbolic VGAE, combining temporal Gnn to model graph dynamics and its uncertainty in hyperbolic space. xu2022uncertainty introduce a Bayesian uncertainty propagation method and embeds Gnns in a Bayesian modeling framework. This approach uses Bayesian confidence of predictive probability based on message uncertainty to quantify predictive uncertainty in node classification. The proposed uncertainty-oriented loss penalizes the predictions with high uncertainty during the learning process, enabling Gnns to improve prediction reliability and OOD detection ability. liu2022ud introduce the concept of a mixture of homophily and heterophily at the node-level, revealing that Gnns may exhibit relatively high model uncertainty for heterophilous nodes. To address this issue, the paper proposes an Uncertainty-aware Debiasing framework, which estimates uncertainty in Bayesian Gnn output to identify heterophilous nodes and then trains a debiased Gnn by pruning biased parameters and retraining the pruned parameters on nodes with high uncertainty.
@@
Edge-level. To better consider the inter-dependencies in the graph structure, some methods also explore uncertainty beyond the node-level. rong2019dropedge introduce DropEdge, which randomly removes a specified number of edges from the graph by drawing independent Bernoulli random variables (with a constant rate) during each training iteration. Serving as both a data augmenter and a message-passing reducer, it alleviates the problems of over-fitting and over-smoothing. hsu2022graph propose two edgewise metrics for graph uncertainty estimation: (1) edgewise expected calibration error (ECE) is designed to estimate the general quality of the predicted edge marginals, utilizing graph structure and considering confidence calibration of edgewise marginals; (2) agree/disagree ECEs extend the idea of edgewise ECE by distinguishing homophilous and heterophilous edges.
@@
Graph-level. Some works address uncertainty in the overall structure or properties of the entire graph. zhang2019bayesian is the first work to develop the Bayesian neural network framework tailored for graphs, treating the observed graph as a realization from a parametric family of random graphs. The joint posterior distribution of the random graph parameters and the node labels are then targeted for inference. However, besides the considerable computation cost, Bayesian Gnn relies on the selection of the random graph model, which may vary across different problems and datasets. hasanzadeh2020bayesian introduce Graph DropConnect (GDC), a general stochastic regularization technique for Gnns through adaptive connection sampling, which yields a mathematically equivalent approximation of Bayesian Gnns training. This unified framework mitigates over-fitting and over-smoothing issues of deep Gnns, while also prompting uncertainty-aware learning. While both Bayesian Gnn and GDC are aware of graph-level uncertainty, the current implementation of posterior inference in Bayesian Gnn solely relies on the graph topology and neglects the node features. In contrast, GDC involves a free parameter that allows adjustment of the binary mask for edges, nodes, and channels, thereby facilitating flexible connection sampling and addressing uncertainty from various levels of graph structure. Therefore, GDC can be viewed as an extension and generalization of DropEdge, which only focuses on edge-level. Moreover, unlike stochastic regularization methods relying on fixed sampling rates or manual hyperparameter tuning (e.g., DropEdge), GDC enables joint training of drop rate and Gnn parameters. liu2020uncertainty propose an Uncertainty-aware Graph Gaussian Process (UaGGP) approach to address uncertainty issues in graph structures, especially when edges connect nodes with different labels, leading to predictive uncertainty. The work employs graph structural information for the feature aggregation and the Laplacian graph regularization. Leveraging the randomness modeled in the Gaussian Process, UaGGP introduces graph Laplacian regularization based on Gaussian Process inferences and symmetric Mahalanobis distance to tackle structural uncertainty.
@@
To address the uncertainty in graph structure, many studies leverage Bayesian inference to capture uncertainty and estimate model parameters. Some approaches can handle over-smoothing and over-fitting phenomena in Gnn simultaneously. In addition, we find these often hold the assumption that edges connecting nodes with different labels would result in misleading neighbors and causing uncertainty in the predictions. Although this assumption is shown valid and effective, we believe it is not the only factor causing uncertainty in graph structure, and this hypothesis is also not suitable for heterogeneous graphs. It would be interesting to design methods that treat structural uncertainty from other geometric information from the graph, e.g., nodes connecting to different communities (huang2023unsupervised, ). Furthermore, while existing works primarily aim to develop uncertainty-aware Gnns to enhance prediction performance, a few explicitly (1) evaluate the reduction of uncertainty in Gnns compared to prior network structures, and (2) validate as well as explain how the reduction of uncertainty contributes to improving prediction accuracy. Finally, the methods in this domain rarely compare against each other as baselines, lacking a universal benchmark.
@@ 4.4 Uncertainty for Trustworthy Gnn
Explainability. Uncertainty quantification provides insights for the reliability and confidence of Gnn decisions. Specifically, mutual information explains the relationship between the Gnns predictions and the components that impact Gnns predictions. ying2019gnnexplainer propose GnnExplainer which identifies a subgraph and/or node features that maximize mutual information with the prediction, to explain which components of the graph have the most impact on the prediction results. Mutual information is nonparametric and does not require prior knowledge and assumptions on models. Thus, it is model-agnostic and suitable for providing interpretable explanations for any Gnn-based models.
Besides, uncertainty quantification can also be used to explain predictions of automated Gnns. yang2022calibrate propose HyperU-GCN, which combines hyperparameter optimization with uncertainty quantification, to explain the selection of hyperparameters. HyperU-GCN quantifies hyperparameter uncertainty as the mutual information between the model prediction and hyperparameters using a probabilistic hypernetwork. HyperU-GCN employs a concept akin to Bayesian Gnns for computing this mutual information, which marginalizes hyperparameters rather than model parameters. Unlike GnnExplainer, HyperU-GCN requires model hyperparameter priors (e.g. normal distribution), yet leveraging Bayesian Gnns enables HyperU-GCN to better calibrate model predictions.
@@
Robustness. Gnns are vulnerable to adversarial attacks, i.e., small perturbations in graph structures and node attributes could lead to performance degradation. Incorporating uncertainty in graph learning empowers Gnns with robustness against such attacks. zhu2019robust propose Robust GCN (RGCN), which adopts Gaussian distributions as the hidden representations of nodes instead of certain vectors. By introducing Gaussian distributions, RGCN can automatically absorb the effects of abnormal adversarial changes and enhance its robustness. Besides, RGCN utilizes a variance-based attention mechanism, which assigns different weights to node neighborhoods according to their variances to reduce the propagated adversarial effects from neighbors. Though RGCN can model epistemic uncertainty in Gnns by Gaussian distribution variances, it does not consider aleatoric uncertainty and uncertainty from graph topology. Thus, feng2021uag propose quantifying uncertainties from graph topology features, measuring both model and data (i.e., aleatoric) uncertainties. This paper utilizes the Bayesian Gnn to estimate model uncertainty and node diversity, which is intrinsic to the graph topology and defined as the number of different labels in the node’s k-hop neighbors, to estimate data uncertainty. Another disadvantage of RGCN is that it relies on prior distributions and the uncertainties are not well-calibrated. Therefore, shanthamallu2021uncertainty propose UM-GNN, which utilizes Bayesian Gnns with Monte Carlo dropout to estimate epistemic uncertainty of a standard Gnn and assigns low uncertainty nodes with high attention. UM-GNN also utilizes an uncertainty-matching strategy to align prediction results between the standard Gnn and a structure-free surrogate model to avoid simple prior assumptions on structural uncertainties. Besides, UM-GNN uses a label smoothing regularization to calibrate uncertainties. zhang2021multi propose a multi-view confidence-calibrated Gnn encoder (MCCG) module, which utilizes Dirichlet distribution to calibrate the prediction confidence for noisy or adversarial samples and utilizes the estimated uncertainty scores to control information propagation. Although these defense methods apply different quantification methods for different sources of uncertainty, they use the same key idea that increasing the weights of low-uncertainty nodes reduces the impact of the adversarial attack.
In conclusion, uncertainty can enhance the trustworthiness of Gnns from both explainability and robustness. Current works reach a consensus on how to utilize uncertainty, e.g., employing mutual information to explain Gnn. Nonetheless, present methods for GNN explainability do not adequately account for the interrelations among components or hyperparameters. Addressing this gap, the application of game-theoretic measures, such as the Shapley value, can be a viable approach. As a common approach in explainable Gnn (kakkad2023survey, ), Shapley value is capable of assessing the individual impact of each component or hyperparameter on the conditional entropy of the model’s output (watson2023explaining, ). Concerning the robustness of Gnns, there are also promising avenues to investigate. For example, instead of increasing the weight of certain nodes, the robustness of Gnn may also be improved by training with perturbations obtained by maximizing the model’s estimated uncertainty (pagliardini2022improving, ).
@@ 5 Uncertainty in Gnns: Real-world Applications
Traffic. Gnns have been widely used to model network traffic, especially traffic demand forecasting. Uncertainty arises in traffic systems both spatially and temporally, e.g., travel demand may increase on game days around the stadiums. wu2021trafficquantifying compare six different Bayesian and Frequentist-based uncertainty quantification methods for road network traffic prediction problems, regarding computational efficiency, asymptotic consistency, etc. zhuang2022trafficuncertainty specifically focus on sparse travel demand and proposes STZINB-GNN, where a spatial-temporal embedding is designed with an additional sparsity parameter to learn the likelihood of inputs being zero. The probability layer in the embedding allows for prediction with a level of confidence and thus, quantifying the demand uncertainty. wang2023trafficuncertainty propos a framework of probabilistic Gnns, Prob-GNN, to quantify spatiotemporal uncertainty of traffic demand. Besides different Gnns, different probabilistic distribution assumptions with different properties are compared, such as Laplace and Gaussian distribution, as well as single and ensembled Gaussian models. Prob-GNN can not only quantify spatiotemporal uncertainty but also generalize well under significant system disruption and identify abnormally large uncertainty. The experiment results show that the Gnn performance is more impacted by probabilistic distribution assumptions than the selection of Gnns, underscoring the importance of incorporating randomness and uncertainty into the traffic prediction model. To quantify both aleatoric and epistemic spatiotemporal uncertainty, qian2023trafficuncertainty propose DeepSTUQ, which uses Bayesian dropout for pre-training and ensembling for re-training, and then calibrates the trained model with temperature scaling. In traffic applications, models are typically evaluated with point estimate error along with the mean prediction interval width (MPIW).
@@
Molecules. Drug discovery is a graph-level problem of identifying and developing new molecules with desired properties, where each molecule is modeled as a graph. zhang2019bayesianchemical compare different Bayesian-based methods for accurate and reliable molecule property predictions in drug discovery. They reveal a few key aspects: (1) both epistemic and aleatoric uncertainty are correlated with error and improve the accuracy of selective prediction, thus, total uncertainty should be utilized for uncertainty estimation; (2) Bayesian uncertainty estimation could overcome the bias of the synthetic dataset by informing the user about the level of uncertainty; (3) the estimated uncertainty is also applied to guide active learning for drug discovery, showing better performance with initially diverse labeled classes. Similarly, ryu2019bayesian apply Bayesian Gnn with Monte Carlo dropout to estimate uncertainty in activity and toxicity classification problems. Additionally, they use aleatoric uncertainty to evaluate the quality of the existing and synthetic datasets. To reduce overconfident mispredictions for drug discovery under distributional shift, han2021reliable introduce CardioTox, a novel real-world OOD data benchmark with molecules annotated by Tanimoto graph distance and proposes a distance-awareness classifier, GNN-GP. Follow-up studies (aouichaoui2022uncertainty, ; kwon2022uncertainty, ) also derive similar conclusions. In molecular force fields, Gnns have also shown promising performance for quantum mechanical calculations to accelerate molecular dynamics trajectory simulations. Besides common concerns about prediction accuracy, uncertainty awareness, and computational costs, wollschlager2023uncertainty propose physics-informed desiderata, including symmetry, energy conservation, and locality. In addition to evaluating previous uncertainty quantification methods, they propose Localized SVGP-DKL, which combines the Bayesian Gnn with the localized neural kernel to satisfy the physics-related criteria.
@@
Others. To segment images of complex geometry or topology, gupta2023topologyaware employ Gnns to jointly reason about the structures (nodes) and capture the high-order spatial interactions (edges). A probabilistic model provides structure-wise uncertainty estimates, which could identify highly uncertain structures for human verification. In computational social systems, zha2023career apply an uncertainty-aware graph encoder (UnGE) to represent job titles as Gaussian embeddings and capture uncertainties in real-world career mobility. In fault diagnosis, a Bayesian hierarchical Gnn is proposed by chen2023bayesian , where the total uncertainty is measured by the dropout. Using uncertainty as a weight, the loss encourages the feature of a highly uncertain sample to be temporally consistent for robust feature learning. Uncertainty in Gnn is also widely applied in diverse domains, including energy (li2023uncertaintyapp, ), robotics (gao2023uncertaintyapp, ), and recommendation systems (zhao2023modeling, ).

### v2

@@ 4.3 Uncertainty-aware Gnn Modeling
Node-level. In graphs, nodes often exhibit conflicting or contradictory characteristics, such as belonging to different communities or revealing contradictory underlying patterns (Bojchevski & Günnemann, 2017). This discrepancy should be reflected in the uncertainty of node embeddings, highlighting the need to account for uncertainty at the node-level in Gnn structures. Hajiramezanali et al. (2019) incorporate stochastic latent variables by combining a Graph Recurrent Neural Network (GRNN) with a Variational Graph Autoencoder (VGAE) (Kipf & Welling, 2016). The proposed variational GRNN is designed to capture the temporal dependencies in dynamic graphs. Moreover, it aims to represent each node by modeling its distribution in the latent space rather than assigning it a deterministic vector in a low-dimensional space, thus enhancing its ability to represent the uncertainty associated with node latent representations. Vashishth et al. (2019) propose ConfGCN, which introduces confidence (inversely proportional to uncertainty) estimation for the label scores on nodes to enhance label predictions. More specifically, ConfGCN assumes a fixed uncertainty for input labels and defines an influence score of each node as a parameter that can be optimized jointly with the network parameters through back-propagation. The influence score here is the inverse of the co-variance matrix-based symmetric Mahalanobis distance between two nodes in the graph. Sun et al. (2021) generate stochastic representations using a hyperbolic VGAE, combining temporal Gnn to model graph dynamics and its uncertainty in hyperbolic space. Xu et al. (2022) introduce a Bayesian uncertainty propagation method and embeds Gnns in a Bayesian modeling framework. This approach uses Bayesian confidence of predictive probability based on message uncertainty to quantify predictive uncertainty in node classification. The proposed uncertainty-oriented loss penalizes the predictions with high uncertainty during the learning process, enabling Gnns to improve prediction reliability and OOD detection ability. Liu et al. (2022e) introduce the concept of a mixture of homophily and heterophily at the node-level, revealing that Gnns may exhibit relatively high epistemic uncertainty for heterophilous nodes. To address this issue, the paper proposes an Uncertainty-aware Debiasing framework, which estimates uncertainty in Bayesian Gnn output to identify heterophilous nodes and then trains a debiased Gnn by pruning biased parameters and retraining the pruned parameters on nodes with high uncertainty.
@@
Edge-level. To better consider the inter-dependencies in the graph structure, some methods also explore uncertainty beyond the node-level. While Veličković et al. (2018) has applied a similar dropout technique on edge attentions, Rong et al. (2019) first formally presented the formulation of DropEdge. It randomly removes a specified number of edges from the graph by drawing independent Bernoulli random variables (with a constant rate) during each training iteration. Rong et al. (2019) theoretically and empirically show that DropEdge, as a data augmenter and a message-passing reducer, can alleviates the problems of over-fitting and over-smoothing. Hsu et al. (2022a) propose two edgewise metrics for graph uncertainty estimation: (1) edgewise expected calibration error (ECE) is designed to estimate the general quality of the predicted edge marginals, utilizing graph structure and considering confidence calibration of edgewise marginals; (2) agree/disagree ECEs extend the idea of edgewise ECE by distinguishing homophilous and heterophilous edges.
@@
Graph-level. Some works address uncertainty in the overall structure or properties of the entire graph. Zhang et al. (2019b) is the first work to develop the Bayesian neural network framework tailored for graphs, treating the observed graph as a realization from a parametric family of random graphs. The joint posterior distribution of the random graph parameters and the node labels are then targeted for inference. However, besides the considerable computation cost, Bayesian Gnn relies on the selection of the random graph model, which may vary across different problems and datasets. Hasanzadeh et al. (2020) introduce Graph DropConnect (GDC), a general stochastic regularization technique for Gnns through adaptive connection sampling, which yields a mathematically equivalent approximation of Bayesian Gnns training. This unified framework mitigates over-fitting and over-smoothing issues of deep Gnns, while also prompting uncertainty-aware learning. While both Bayesian Gnn and GDC are aware of graph-level uncertainty, the current implementation of posterior inference in Bayesian Gnn solely relies on the graph topology and neglects the node features. In contrast, GDC involves a free parameter that allows adjustment of the binary mask for edges, nodes, and channels, thereby facilitating flexible connection sampling and addressing uncertainty from various levels of graph structure. Therefore, GDC can be viewed as an extension and generalization of DropEdge, which only focuses on edge-level. Moreover, unlike stochastic regularization methods relying on fixed sampling rates or manual hyperparameter tuning (e.g., DropEdge), GDC enables joint training of drop rate and Gnn parameters. Liu et al. (2020) propose an Uncertainty-aware Graph Gaussian Process (UaGGP) approach to address uncertainty issues in graph structures, especially when edges connect nodes with different labels, leading to predictive uncertainty. The work employs graph structural information for the feature aggregation and the Laplacian graph regularization. Leveraging the randomness modeled in the Gaussian Process, UaGGP introduces graph Laplacian regularization based on Gaussian Process inferences and symmetric Mahalanobis distance to tackle structural uncertainty.
@@
To address the uncertainty in graph structure, many studies leverage Bayesian inference to capture uncertainty and estimate model parameters. Some approaches can handle over-smoothing and over-fitting phenomena in Gnn simultaneously. In addition, we find these often hold the assumption that edges connecting nodes with different labels would result in misleading neighbors and causing uncertainty in the predictions. Although this assumption is shown valid and effective, we believe it is not the only factor causing uncertainty in graph structure, and this hypothesis is also not suitable for heterogeneous graphs. It would be interesting to design methods that treat structural uncertainty from other geometric information from the graph, e.g., nodes connecting to different communities (Huang et al., 2023b). Furthermore, while existing works primarily aim to develop uncertainty-aware Gnns to enhance prediction performance, a few explicitly (1) evaluate the reduction of uncertainty in Gnns compared to prior network structures, and (2) validate as well as explain how the reduction of uncertainty contributes to improving prediction accuracy. Finally, the methods in this domain rarely compare against each other as baselines, lacking a universal benchmark.
@@ 4.4 Uncertainty for Trustworthy Gnn
Explainability. Uncertainty quantification provides insights for the reliability and confidence of Gnn decisions. Specifically, mutual information explains the relationship between the Gnns predictions and the components that impact Gnns predictions. Ying et al. (2019) propose GnnExplainer which identifies a subgraph and/or node features that maximize mutual information with the prediction, to explain which components of the graph have the most impact on the prediction results. Mutual information is nonparametric and does not require prior knowledge and assumptions on models. Thus, it is model-agnostic and suitable for providing interpretable explanations for any Gnn-based models. Besides, uncertainty quantification can also be used to explain predictions of automated Gnns. Yang et al. (2022) propose HyperU-GCN, which combines hyperparameter optimization with uncertainty quantification, to explain the selection of hyperparameters. It quantifies hyperparameter uncertainty as the mutual information between the model prediction and hyperparameters using a probabilistic hypernetwork. HyperU-GCN employs a concept akin to Bayesian Gnns for computing this mutual information, which marginalizes hyperparameters rather than model parameters. Unlike GnnExplainer, HyperU-GCN requires model hyperparameter priors (e.g. normal distribution), yet leveraging Bayesian Gnns enables it to better calibrate model predictions. There are many other methods to interpret Gnns besides using uncertainty, and we refer readers to Kakkad et al. (2023) for a comprehensive overview. For example, GraphLIME (Huang et al., 2022) samples neighbors and finds the most representative features as Gnn local explanations based on Hilbert-Schmidt Independence Criterion Lasso; GCExplainer (Magister et al., 2021) applies an Automated Concept-based Explanation approach (Ghorbani et al., 2019), which maps node representations to the concept space, to extract important concepts as global Gnn explanations.
@@
Robustness. Gnns are vulnerable to adversarial attacks, i.e., small perturbations in graph structures and node attributes could lead to performance degradation. Incorporating uncertainty in graph learning empowers Gnns with robustness against such attacks. Zhu et al. (2019) propose Robust GCN (RGCN), which adopts Gaussian distributions as the hidden representations of nodes instead of certain vectors. By introducing Gaussian distributions, RGCN can automatically absorb the effects of abnormal adversarial changes and enhance its robustness. Besides, RGCN utilizes a variance-based attention mechanism, which assigns different weights to node neighborhoods according to their variances to reduce the propagated adversarial effects from neighbors. Though RGCN can model epistemic uncertainty in Gnns by Gaussian distribution variances, it does not consider aleatoric uncertainty and uncertainty from graph topology. Thus, Feng et al. (2021) propose quantifying uncertainties from graph topology features, measuring both epistemic and aleatoric (i.e., data) uncertainties. This paper utilizes the Bayesian Gnn to estimate epistemic uncertainty and node diversity, which is intrinsic to the graph topology and defined as the number of different labels in the node’s k-hop neighbors, to estimate aleatoric uncertainty. Another disadvantage of RGCN is that it relies on prior distributions and the uncertainties are not well-calibrated. Therefore, Shanthamallu et al. (2021) propose UM-GNN, which utilizes Bayesian Gnns with Monte Carlo dropout to estimate epistemic uncertainty of a standard Gnn and assigns low uncertainty nodes with high attention. UM-GNN also utilizes an uncertainty-matching strategy to align prediction results between the standard Gnn and a structure-free surrogate model to avoid simple prior assumptions on structural uncertainties. Besides, UM-GNN uses a label smoothing regularization to calibrate uncertainties. Zhang et al. (2021c) propose a multi-view confidence-calibrated Gnn encoder (MCCG) module, which utilizes Dirichlet distribution to calibrate the prediction confidence for noisy or adversarial samples and utilizes the estimated uncertainty scores to control information propagation. Although these defense methods apply different quantification methods for different sources of uncertainty, they use the same key idea that increasing the weights of low-uncertainty nodes reduces the impact of the adversarial attack. Besides, although it is not explicitly stated that uncertainty quantification methods were used, some works have employed uncertainty-related approaches to verify the certifiable robustness of Gnns. For example, Bojchevski & Günnemann (2019) uses the minimum worst-case margin between the predicted or ground-truth node label and other labels under perturbations of graph structures to verify whether the node is robust. By intercepting messages from perturbed nodes, Scholten et al. (2022) uses whether the perturbed prediction is consistent with its neighbors regarding the majority vote to certify the robustness of Gnns. Both works perturb the original input graphs and measure the robustness with the variability of predictions under perturbations. This idea is similar to Bayesian Gnns, where uncertainty is measured with the variance of prediction with randomized Gnn parameters. In this way, robust Gnns are expected to preserve consistent and certain predictions with noisy data. Sun et al. (2022) provides more Gnns robustness studies.
In conclusion, uncertainty can enhance the trustworthiness of Gnns from both explainability and robustness, with a consensus on how to utilize uncertainty. Other interesting directions are also worth exploring. For example, instead of increasing the weight of certain nodes, the robustness of Gnn may also be improved by training with perturbations obtained by maximizing the model’s estimated uncertainty (Pagliardini et al., 2022).
@@ 5 Uncertainty in Gnns: Real-world Applications
Traffic. Gnns have been widely used to model network traffic, especially traffic demand forecasting. Uncertainty arises in traffic systems both spatially and temporally, e.g., travel demand may increase on game days around the stadiums. Wu et al. (2021) compare six different Bayesian and Frequentist-based uncertainty quantification methods for road network traffic prediction problems, regarding computational efficiency, asymptotic consistency, etc. Zhuang et al. (2022) specifically focus on sparse travel demand and proposes STZINB-GNN, where a spatial-temporal embedding is designed with an additional sparsity parameter to learn the likelihood of inputs being zero. The probability layer in the embedding allows for prediction with a level of confidence and thus, quantifying the demand uncertainty. Wang et al. (2023) propos a framework of probabilistic Gnns, Prob-GNN, to quantify spatiotemporal uncertainty of traffic demand. Besides different Gnns, different probabilistic distribution assumptions with different properties are compared, such as Laplace and Gaussian distribution, as well as single and ensembled Gaussian models. Prob-GNN can not only quantify spatiotemporal uncertainty but also generalize well under significant system disruption and identify abnormally large uncertainty. The experiment results show that the Gnn performance is more impacted by probabilistic distribution assumptions than the selection of Gnns, underscoring the importance of incorporating randomness and uncertainty into the traffic prediction model. To quantify both aleatoric and epistemic spatiotemporal uncertainty, Qian et al. (2023) propose DeepSTUQ, which uses Bayesian dropout for pre-training and ensembling for re-training, and then calibrates the trained model with temperature scaling. In traffic applications, models are typically evaluated with point estimate error along with the mean prediction interval width (MPIW).
@@
Molecules. Drug discovery is a graph-level problem of identifying and developing new molecules with desired properties, where each molecule is modeled as a graph. Zhang et al. (2019a) compare different Bayesian-based methods for accurate and reliable molecule property predictions in drug discovery. They reveal a few key aspects: (1) both epistemic and aleatoric uncertainty are correlated with error and improve the accuracy of selective prediction, thus, total uncertainty should be utilized for uncertainty estimation; (2) Bayesian uncertainty estimation could overcome the bias of the synthetic dataset by informing the user about the level of uncertainty; (3) the estimated uncertainty is also applied to guide active learning for drug discovery, showing better performance with initially diverse labeled classes. Similarly, Ryu et al. (2019) apply Bayesian Gnn with Monte Carlo dropout to estimate uncertainty in activity and toxicity classification problems. Additionally, they use aleatoric uncertainty to evaluate the quality of the existing and synthetic datasets. To reduce overconfident mispredictions for drug discovery under distributional shift, Han et al. (2021) introduce CardioTox, a novel real-world OOD data benchmark with molecules annotated by Tanimoto graph distance and proposes a distance-awareness classifier, GNN-GP. Follow-up studies (Aouichaoui et al., 2022; Kwon et al., 2022) also derive similar conclusions. In molecular force fields, Gnns have also shown promising performance for quantum mechanical calculations to accelerate molecular dynamics trajectory simulations. Besides common concerns about prediction accuracy, uncertainty awareness, and computational costs, Wollschläger et al. (2023) propose physics-informed desiderata, including symmetry, energy conservation, and locality. In addition to evaluating previous uncertainty quantification methods, they propose Localized SVGP-DKL, which combines the Bayesian Gnn with the localized neural kernel to satisfy the physics-related criteria.
@@
Others. To segment images of complex geometry or topology, Gupta et al. (2023) employ Gnns to jointly reason about the structures (nodes) and capture the high-order spatial interactions (edges). A probabilistic model provides structure-wise uncertainty estimates, which could identify highly uncertain structures for human verification. In computational social systems, Zha et al. (2023) apply an uncertainty-aware graph encoder (UnGE) to represent job titles as Gaussian embeddings and capture uncertainties in real-world career mobility. In fault diagnosis, a Bayesian hierarchical Gnn is proposed by Chen et al. (2023), where the total uncertainty is measured by the dropout. Using uncertainty as a weight, the loss encourages the feature of a highly uncertain sample to be temporally consistent for robust feature learning. Uncertainty in Gnn is also widely applied in diverse domains, including energy (Li et al., 2023a), robotics (Gao et al., 2023), and recommendation systems (Zhao et al., 2023).